OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x − 1) less than or greater to 10? x 6 6 x A number line with a closed circle on 6 and shading to the left. A number line with a closed circle on 6 and shading to the right.

OpenStudy (anonymous):

\(2(x-1) \leq 10\) Have you solved the inequality for x?

OpenStudy (anonymous):

correction x <= 6 6 => X

OpenStudy (anonymous):

yes I have its X <= 6

OpenStudy (anonymous):

OK. The "or equal" part means that 6 is included and that is indicated with a closed circle on the number line.

OpenStudy (anonymous):

The shaded area are all the numbers that could be true for this equation.

OpenStudy (anonymous):

i know there are conflicting answers A and D are the same

OpenStudy (anonymous):

I'm not sure what your answer options actually look like. I'm not sure what x 6 and 6 x are supposed to be. I can help you draw the solution on a number line.

OpenStudy (anonymous):

its one of these A number line with a closed circle on 6 and shading to the left 6 => x which one is it

OpenStudy (anonymous):

In which direction are numbers less than or equal to 6?

OpenStudy (anonymous):

Based on the wording of the question, there may be more than one answer.

OpenStudy (anonymous):

i know answer is x = 6 which is not representative of that A number line with a closed circle on 6 and shading to the left. A number line with a closed circle on 6 and shading to the right. x>=6

OpenStudy (anonymous):

\(x \leq 6\) is represented on the numberline with a closed circle (because it is equal to) and shading to the left to indicate all the numbers less than 6. Any numbers greater than 6 fall to the right and do not solve the inequality.

OpenStudy (anonymous):

it seems there are 2 answers when there should only be 1.

OpenStudy (anonymous):

What are the first two options again?

OpenStudy (anonymous):

x <= 6 6 =>

OpenStudy (anonymous):

Those both say the same thing.

OpenStudy (anonymous):

There is only one answer to this question.

OpenStudy (anonymous):

i know this is what the ansewers are A number line with a closed circle on 6 and shading to the left. A number line with a closed circle on 6 and shading to the right x <= 6 6=> X

OpenStudy (anonymous):

you are looking for the answer that is incorrect. \(x \leq 6\) and \(6 \geq x\) say the same thing as the numberline with a closed circle on the 6 and shading to the left.

OpenStudy (anonymous):

ok i have a couple more questions ] Which of the following is not a way to represent the solution of the inequality 3(2x − 1) greater than or equal to 4(2x − 3) − 3? x less than or greater to 6 6 greater than or equal to x A number line with a closed circle on 6 and shading to the left. A number line with a closed circle on 6 and shading to the right

OpenStudy (anonymous):

\[3(2x-1) \geq 4(2x-3)-3\] Did you find the solution?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

The thing to remember here is that when you multiply or divide by a negative, you have to reverse the inequality.

OpenStudy (anonymous):